OpenStudy (anonymous):

a steel wire of length 4.7 m and cross section 3*10^-5 m^2 stretches by the same amount as a copper wire of length 3.5 m and cross section 4 * 10^-5 m^2 . Find the ratio of the youngs modulus of steel to that of copper. @IIT study group

OpenStudy (anonymous):

well I don't study about Youngs modulus, but I guess you can find the answer by using this formula. \[ E \equiv \frac{\mbox {tensile stress}}{\mbox {tensile strain}} = \frac{\sigma}{\varepsilon}= \frac{F/A_0}{\Delta L/L_0} = \frac{F L_0} {A_0 \Delta L} \] F is the force exerted on an object under tension ΔL is the amount by which the length of the object changes A0 is the original cross-sectional area through which the force is applied L0 is the original length of the object (for more details, go Wiki) Anyway, I think the force 'F' and change of the object ΔL is same, so E(S-steel)/E(C-copper)= A0(c)*L0(s)/A(0)(s)*L0(c) Therefore, \[[4\times10^{-5}][4.7]/[3\times10^{-5}][3.5]=18.8/10.5=1.8\]

OpenStudy (anonymous):

sorry...cant see the answer part..is the answer 18.8/10.5? thank u..

OpenStudy (anonymous):

can u please help me in 1 more question??